# LLM-RankFusion: Mitigating Intrinsic Inconsistency in LLM-based Ranking

### v1

@@ 5.3 Addressing Inconsistency via Aggregation
Table 5: NDCG@10 results on aggregating across sorting algorithms. We show the standard deviation for each NDCG@10 value among 100 experiments.
w/o ICL + Calibration w/ ICL + Calibration LLM-RankFusion
Bubblesort Heapsort Bubblesort Heapsort Aggregated
Flan-T5-XXL 0.377 0.202 0.190 0.075 0.115
Flan-UL2 0.359 0.211 0.165 0.069 0.104
Llama-3-8B 0.408 0.197 0.130 0.060 0.084
Llama-3-70B 0.352 0.228 0.147 0.066 0.095
Gemma-7B 0.493 0.430 0.187 0.075 0.115
Vicuna-13B 0.444 0.272 0.151 0.067 0.096
Mixtral-8x7B 0.449 0.290 0.299 0.123 0.192
GPT-3.5-Turbo 0.434 0.242 0.172 0.076 0.108
GPT-4 0.332 0.163 0.229 0.142 0.173
Table 6: Average Kendall-tau distance results in aggregating across sorting algorithms. The standard deviation for Kendall-tau distance is too small and not significant, so we don’t include it in this table.
Aggregation across sorting algorithms
attempts to mitigate the impact of the initial order on an individual sorting algorithm. As seen in Table 5 and Table 6, the LLM-RankFusion pipeline provides significant benefits to baseline sorting-based pairwise approaches. Each of the initial orderings underwent re-ranking from both pairwise prompting with Bubblesort and Heapsort, and the two final rank lists were aggregated using Borda count. NDCG@10 and Kendall-tau distance were recorded and averaged across each final list for each model. In all models, we see an increase in NDCG@10 as well as reduced inconsistency with lower Kendall-tau distances. ICL and calibration significantly increase the NDCG while also reducing the Kendall-tau distance, and the aggregation further reduces the inconsistency of the worse sorting algorithm (in this case Bubblesort), while keeping NDCG relatively constant. This is a promising result, showing that LLM-RankFusion can improve both consistency and ranking quality by aggregating the rankings produced by different sorting algorithms.
Bubblesort Heapsort
KT NDCG@10 KT NDCG@10
Small-Model LLM-RankFusion 0.187 69.50 0.062 71.32
Medium-Model LLM-RankFusion 0.226 70.93 0.076 73.09
Large-Model LLM-RankFusion 0.197 74.22 0.084 74.75
Table 7: Aggregation results across 100 initial orderings for multiple models based on respective size. Small-model aggregation consists of Flan-T5-XXL, Gemma-7b, and Llama-3-8B. Medium-Model Aggregation contains Flan-UL2, Vicuna-13b, and Mixtral-8x7b. Large-model aggregation reflects aggregation of Llama-3-70b, GPT-3.5-Turbo, and GPT-4. The fully detailed version of this table is shown in Table 8.
Aggregation across LLMs
attempts to produce a final ranking that is less sensitive to the individual biases of different LLMs. This can be observed by comparing Table 7 to previous Table 5 and Table 6, as the average Kendall-tau distance of the aggregated rankings can always achieve a medium performance among the LLMs it aggregates from. This shows promising results of LLM-RankFusion to find a consensus ranking list that can achieve medium performance. In a typical passage ranking case, we don’t assume we know which LLM is better at ranking the specific passage list. Hence, it is particularly useful if we can get the medium performance rankings by aggregating from multiple proposals.
5.4 Computation Cost
The performance of LLM-based passage ranking is mainly determined by the LLM inference performance and the number of comparisons required to sort a list. The aggregation in LLM-RankFusion requires that all the individual ranking lists be fully sorted, which means the time cost will be the sum of the individual ranking time. The calibration doesn’t bring additional inference costs because it only collects the output logits and calculates preference probability based on two permutations of passages. The ICL leads to a longer prompt, which increases the computation during inference, as shown in Table 13. The performance of ranking individual passage lists is the same as previous PRP-Sorting [32] for those without ICL cases.
6 Conclusion
Limitations
The aggregation process in LLM-RankFusion requires fully sorted individual ranked lists, which increases the computation costs. Additionally, the use of ICL leads to longer prompts, increasing the overhead during inference. Another limitation of our study is that the experiments were conducted on a specific dataset (TREC-DL2019), and other aggregation methods were not included in the current research.
Future Work
can focus on LLM-based rank aggregation approaches that decide the comparison strategy on-the-fly and directly aggregate from pairwise comparisons without relying on sorting algorithms [33, 16, 15]. Exploring the potential of aggregating LLM-based decisions in other tasks and domains beyond passage ranking could lead to a more general understanding of the effectiveness of combining the intelligence of multiple LLMs for improved performance and consistency in a wider range of applications [11].
Appendix A Appendix
$A$$B$$C$$D$
CDBACDABBADCABCDBubblesortBubblesortInitial OrderFinal Order
Figure 2: Effects of LLM inconsistency on the results of sorting-based reranking. The graph on the left shows the LLM’s pairwise comparisons of documents $A,B,C,D$, which have multiple transitive inconsistencies. A directed edge from A to B represents A being preferred over B. When using these pairwise comparisons to perform a Bubblesort of the four documents, the final sorted order can vary drastically based on initial ordering, as shown on the right.
A.1 Additional Experiments
We compare the LLM-RankFusion on aggregating among multiple models in Table 8, which is also a detailed version of Table 7. In table 8, we illustrate that the multi-model aggregation also balances out worse NDCG@10 and Kendall-tau distances associated with specific models, and provides a more comprehensive ranking result taking multiple factors into account. The bubblesort and heapsort data associated with each model refers to after ICL and calibration have been applied, so this table directly looks at the impact of aggregation specifically.
@@ A.1 Additional Experiments
Table 8: Aggregation results across 100 initial orderings for multiple models based on respective size. All results use ICL + Calibration.
A.2 Example Passage Ranking

### v2

@@ 5.3 Addressing Inconsistency via Aggregation
Table 6 demonstrates the effectiveness of both model and ranking scheme aggregation in improving ranking performance. Model Aggregation consistently outperforms individual models across both datasets, particularly for pairwise ranking. This trend is consistent across different ranking schemes and datasets, showing robustness through aggregation. On ranking scheme aggregation, we observe that combining different approaches mitigates individual weaknesses. Scheme Aggregation shows balanced performance, often achieving scores close to or exceeding the best individual scheme. This suggests that aggregating across ranking schemes can provide a more stable and potentially superior ranking, leveraging the strengths of both pairwise and setwise approaches, as well as different sorting algorithms.
@@
Table 7: NDCG@10 of aggregating across sorting algorithms with standard deviation among 100 experiments on TREC DL 2019.
@@
Table 8: Aggregation results across 100 initial orderings for multiple models based on respective size on TREC DL 2019. It shows that the multi-model aggregation also balances out worse NDCG@10 and Kendall-tau distances associated with specific models, and provides a more comprehensive ranking result taking multiple factors into account. The bubblesort and heapsort data associated with each model refer to after ICL and calibration have been applied.
Aggregation across sorting algorithms
attempts to mitigate the impact of the initial order on an individual sorting algorithm. As seen in Table 7 and Table 9 in the appendix, LLM-RankFusion provides significant benefits to baseline sorting-based ranking approaches. Each of the initial orderings underwent re-ranking from both pairwise prompting with Bubblesort and Heapsort, and the two final rank lists were aggregated using Borda count. NDCG@10 and Kendall-tau distance were recorded and averaged across each final list for each model. In all models, we see an increase in NDCG@10 as well as reduced inconsistency with lower Kendall-tau distances. ICL and calibration significantly increase the NDCG while also reducing the Kendall-tau distance, and the aggregation further reduces the inconsistency of the worse sorting algorithm (in this case Bubblesort), while keeping NDCG relatively constant. This is a promising result, showing that LLM-RankFusion can improve both consistency and ranking quality by aggregating the rankings produced by different sorting algorithms.
Aggregation across LLMs
attempts to produce a final ranking that is less sensitive to the individual biases of different LLMs. This can be observed by comparing Table 8 to previous Table 7 and Table 9 in the appendix, as the average Kendall-tau distance of the aggregated rankings can always achieve a medium performance among the LLMs it aggregates from. This shows promising results of LLM-RankFusion to find a consensus ranking list that achieves balanced or even superior performance. In a typical passage ranking case, we don’t assume we know which LLM is better at ranking the specific passage list. Hence, it is particularly useful if we can get at least medium performance rankings by aggregating multiple proposals.
5.4 Computation Cost
The performance of LLM-based passage ranking is mainly determined by the LLM inference performance and the number of comparisons required to sort a list. The aggregation in LLM-RankFusion requires that takes in individual ranking lists and the time cost depends on the slowest model if individual rankings are computed in parallel.The calibration doesn’t bring additional inference costs because it only collects the output logits and calculates preference probability based on two permutations of passages. The ICL leads to a longer prompt, which slightly increases the computation during inference as shown in Table 14 in the appendix.
6 Conclusion
Appendix A Technical Appendix
A.1 Limitations
The aggregation process in LLM-RankFusion requires fully sorted individual ranked lists, which increases the computation costs. Additionally, the use of ICL leads to longer prompts, increasing the overhead during inference. Another limitation of our study is that the experiments were conducted on limited IR datasets (TREC-DL2019 and TREC-DL2020), and other aggregation methods were not included in the current research.
A.2 Future Work
Future Work can focus on LLM-based rank aggregation approaches that decide the comparison strategy on-the-fly and directly aggregate from pairwise comparisons without relying on sorting algorithms [33, 16, 15]. Exploring the potential of aggregating LLM-based decisions in other tasks and domains beyond passage ranking could lead to a more general understanding of the effectiveness of combining the intelligence of multiple LLMs for improved performance and consistency in a wider range of applications [11].
w/o ICL + Calibration w/ ICL + Calibration LLM-RankFusion
Bubblesort Heapsort Bubblesort Heapsort Aggregated
Flan-T5-XXL 0.377 0.202 0.190 0.075 0.115
Flan-UL2 0.359 0.211 0.165 0.069 0.104
Llama-3-8B 0.408 0.197 0.130 0.060 0.084
Llama-3-70B 0.352 0.228 0.147 0.066 0.095
Gemma-7B 0.493 0.430 0.187 0.075 0.115
Vicuna-13B 0.444 0.272 0.151 0.067 0.096
Mixtral-8x7B 0.449 0.290 0.299 0.123 0.192
GPT-3.5-Turbo 0.434 0.242 0.172 0.076 0.108
GPT-4 0.332 0.163 0.229 0.142 0.173
Table 9: Average Kendall-tau distance results in aggregating across sorting algorithms. We don’t include the standard deviation for Kendall-tau distance in this table because it is too small and insignificant.
A.3 Example Passage Ranking
@@ A.6 Performance
A.7 Analysis of Prompt Design Sensitivity
Model Baseline ICL Only Calibration Only ICL + Calibration
Bubblesort 62.68±3.07 67.31±1.73 69.59±1.89 70.84±1.61
Heapsort 63.82±4.63 69.45±1.92 68.30±2.24 70.10±1.92
Table 15: NDCG@10 scores (mean ± std) across different prompt designs
To evaluate the robustness of our proposed methods against variations in prompt design, we conducted additional experiments using GPT-4 to generate 10 different pairwise ranking prompt templates. Each template maintained the core task of comparing passage relevance while varying factors such as wording, formatting, and instruction style. We tested these prompts using Llama-3-8B on the TREC-DL2019 dataset under different configurations. The results in Table 15 demonstrate two key benefits of our approach:
  1. 1.
    Enhanced Performance: The combination of ICL and calibration substantially improves ranking quality, with NDCG scores increasing by approximately 8 points compared to the baseline across both sorting algorithms.
  2. 2.
    Reduced Variance: The baseline exhibits considerable sensitivity to prompt design, as evidenced by the high standard deviations (3.07-4.63). In contrast, configurations using ICL and calibration show markedly lower variance (1.61-1.92), indicating more stable performance across different prompt designs.